Solve the system by addition.
–3x + 5y = –18
x – 2y = 7

Answer by checkley75(3666)

-3X+5Y=-18 NOW MULTIPLY THE SECOND EQUATION BY 3 & ADD:
3X-6Y=21 NOW ADD
-----------------------------------
-Y=3
Y=-3 ANSWER.
-3X+5*-3=-18
-3X-15=-18
-3X=-18+15
-3X=-3
X=-3/-3
X=1 ANSWER.
PROOF
1-2*-3=7
1+6=7
7=7